The area of a square is 64 square inches. What is the length of each side?
Answer by edjones(7311)

A=Area, s=side, sqrt=square root
s^2=A
s=sqrt(A) Take the sqrt of each side of the equation.
s=sqrt(64)
s=8 inches
.
Ed